OpenStudy (anonymous):

"3 is what percent of sixty-five?

OpenStudy (anonymous):

3 is what percent of sixty-five 3 = x /100 * 65 3 = x/100 * 65 Solve for x

OpenStudy (anonymous):

\[\frac{ 3 }{ 65 }\] = 4.615%